A random sample of 27 fields of rye has a mean yield of 43.1 bushels per acre and standard deviation of 5.31 bushels per acre. Determine the 80% confidence interval for the true mean yield. Assume the population is approximately normal.

Step 1 of 2 : Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places
Answer #1

Solution :

Given that,

\bar x = 43.1

s = 5.31

n = 27

Degrees of freedom = df = n - 1 = 27 - 1 = 26

At 80% confidence level the t is ,

\alpha = 1 - 80% = 1 - 0.80 = 0.20

\alpha / 2 = 0.20 / 2 = 0.10

t\alpha /2,df = t0.10,26 = 1.315

Critical value = 1.315

Margin of error = E = t\alpha/2 * (s /\sqrtn) = 1.315 * (5.31 / \sqrt27) = 1.34

The 80% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

43.1 - 1.34 < \mu < 43.1 + 1.34

41.76 < \mu < 44.44

(41.76 , 44.44)